in the formula, c represents the total charge in dollars for babysitting, and h represents the number of hours the child is watched. how much should jonathan pay id his child is sat the babysitting service for 3 hours?
c = 5.25 + 4.25 H

Answer by Earlsdon(6294)

c = 5.25 + 4.25H Substitute H = 3 for 3 hours of baby-sitting.
c = 5.25 + 4.25(3)
c = 5.25 + 12.75
c = 18
Jonathan should pay $18.00